OpenStudy (anonymous):

integral of (sinx-cosx)/√(cos^4(x) +sin^4(x))

OpenStudy (anonymous):

Is this your function \[ \frac{\sin (x)-\cos (x)}{\sqrt{\sin ^4(x)+\cos ^4(x)}} \]

OpenStudy (anonymous):

Mathematica could not do it in a closed form. I will be interested if someone can get it.